# Entanglement suppression and quantum scars in a three-oscillator gravitational analogue

### v1

@@ IV Numerical Analysis and physical regimes
To move beyond the analytically tractable weak-coupling regime, we numerically simulate the system’s dynamics using the full Bogoliubov coefficients ($\bm{\alpha},\bm{\beta}$) (See (B) for details). As mentioned above, our analysis focuses on two distinct physical regimes characterized by the mediator’s behavior, considering both initial squeezed and vacuum states. We present here the results for squeezed states, see Appendix (E) for vacuum initial state. The two limits delineate how the mediator’s dynamical accessibility and quantumness control entanglement generation across the system.
@@
    HMR: This regime is defined by the limit $\lambda_{B}\to\omega_{B}$, with $1-\lambda_{A}/\omega_{A}$ remaining finite. In this limit, the mediator oscillator B behaves as an effectively heavy, quasi-classical field mode. To understand its physical implications, consider the single-oscillator Hamiltonian with self-squeezing (cf. Eq. (2b)). When expressed in the position–momentum basis, after transformation, the Hamiltonian simplifies to:
    $\hat{H}_{0B}={\hat{p}_{B}^{2}}/{(2\tilde{m}_{B}})+\tilde{k}_{B}\,\hat{x}_{B}^{2}/2$ (10)
    Here, $\tilde{m}_{B}$ and $\tilde{k}_{B}$ are the renormalized parameters, given by:
    $\tilde{m}_{B}=\frac{m_{B}}{1-\lambda_{B}/\omega_{B}},~{}\tilde{\omega}_{B}^{2}=\omega_{B}^{2}-\lambda_{B}^{2},~{}\tilde{k}_{B}=\tilde{m}_{B}\tilde{\omega}_{B}^{2},$ (11)
    where $m_{B}$ is the bare mass of the mediator. In the critical limiting case $\lambda_{B}\to\omega_{B}$, the effective mass $\tilde{m}_{B}$ diverges ($\tilde{m}_{B}\to\infty$) while the renormalized frequency vanishes ($\tilde{\omega}_{B}\to 0$). Crucially, their product $\tilde{k}_{B}=\tilde{m}_{B}\tilde{\omega}_{B}^{2}$ remains finite and equal to twice the bare spring constant $m_{B}\omega_{B}^{2}$. Physically, this causes the kinetic term to vanish from the Hamiltonian (10), i.e.,
    $\hat{H}_{0B}\xrightarrow[\lambda_{B}\to\omega_{B}]{}\frac{1}{2}\tilde{k}_{B}\,\hat{x}_{B}^{2}.$ (12)
    This limit corresponds to an extreme self-interaction that effectively suppresses the canonical momentum $\hat{p}_{B}$, making $\hat{x}_{B}$ a conserved quantity. Hence, the oscillator no longer evolves dynamically but acts as a static background field fixed by its initial position $\hat{x}_{B}(0)$. This behavior is analogous to a cosmological constant [47], where vacuum energy is given by the constant scalar field potential, here represented by a fixed energy density parameterized by $\hat{x}_{B}(0)^{2}$.
    Thus, in HMR, the mediator loses its ability to coherently store or transmit quantum information. Its Hilbert space remains intact, but its dynamical role is effectively frozen: the canonical momentum $\hat{p}_{B}$ vanishes from the Hamiltonian, and the position $\hat{x}_{B}$ becomes a conserved quantity. This reduction mirrors constrained dynamics in gauge theories and systems with Hamiltonian constraints, where certain degrees of freedom become dynamically redundant or decouple from the evolution [48, 49].
    To visualize how entanglement between subsystems A and C depend on system parameters in both mediator regimes, we plot contour maps of the long-time averaged $\langle\mathcal{E}_{\mathcal{N}}\rangle$. (For $\langle\mathcal{D}\rangle$, see Appendix (E).) The first two plots on the left in Fig. (2) contain the plot of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and $\langle\mathcal{F}\rangle$ respectively as a function of $\omega_{A}$ and $g_{A}$ by fixing $\omega_{B}$, $\lambda_{A}$, $\lambda_{B}$, and $g_{B}$.
@@
    LMR: This regime is characterized by $\lambda_{A}\to\omega_{A}$, while $1-\lambda_{B}/\omega_{B}$ remains finite. In this scenario, the mediator B is not significantly self-squeezed, remaining highly quantum and susceptible to fluctuations. The last two plots on the right in Fig. (2) contain the plot of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and $\langle\mathcal{F}\rangle$ respectively as a function of $\omega_{B}$ and $g_{B}$ by fixing $\omega_{A}$, $\lambda_{A}$, $\lambda_{B}$, and $g_{A}$. (For $\langle\mathcal{D}\rangle$, see Appendix (E).)
@@ V Results and Implications
Our numerical analysis, presented in Fig. 2 reveal that both (HMR and LMR) regimes can support generation of quantum correlations, quantified by $\langle\mathcal{E}_{\mathcal{N}}\rangle$ and $\langle\mathcal{D}\rangle$, between subsystems A and C. However, these regimes exhibit markedly different structural and spectral features in their entanglement.
The LMR consistently exhibits a smoother, more continuous entanglement landscape. Here, non-zero quantum correlations are broadly distributed across wider parameter ranges, indicating a robust ability of the “light” (dynamically active) mediator B to facilitate entanglement between A and C. This suggests that when the mediator can fluctuate significantly, it readily mediates quantum correlations.
In stark contrast, HMR presents a highly fragmented entanglement landscape. Entanglement emerges only within sharply localized islands in parameter space, surrounded by expansive regions where quantum correlations are strongly suppressed, often vanishing entirely. Despite the system’s integrability, the emergence of such sharply structured regions in the HMR is strikingly reminiscent of quantum many-body scars. In systems with scars, specific non-generic eigenstates exhibit unexpectedly low entanglement and resist thermalization, maintaining enhanced localization in Hilbert space. Our findings suggest an analogous form of scars, not necessarily in the many-body Hilbert space, but within the space of Gaussian states of this continuous-variable system. Further fidelity analysis support this: in the low-entanglement regions, the fidelity of the evolved state with the initial state remains remarkably close to unity over long times. This indicates dynamical stability and a resistance to spreading in the Hilbert space, akin to localized scarred states. Conversely, within the high-entanglement islands, fidelity oscillates significantly away from $1$, reflecting more complex, delocalized, and dynamically richer behavior. This coexistence of dynamically localized and delocalized states within an integrable system points towards selective resonance effects or subtle hidden symmetries that prevent full ergodic mixing in the continuous-variable phase space.
These differences are greatly reflected in the underlying eigenmode structure of the system, characterized by the normal mode frequencies $k_{1},k_{2},k_{3}$. In the HMR ($\lambda_{B}\to\omega_{B}$), one eigenmode approaches zero ($k_{2}\to 0$), signifying the “freezing” of the mediator’s kinetic degree of freedom, while the other two become degenerate ($k_{1}\to k_{3}$), implying a symmetric, coupled motion of A and C dominated by the static mediator. In the LMR ($\lambda_{A}\to\omega_{A}$), a different spectral signature emerges: two eigenmodes vanish simultaneously ($k_{1},k_{2}\to 0$), with the third remaining finite. This indicates a different type of effective decoupling or simplification of dynamics, possibly leaving the mediator’s own quantum fluctuations more dominant in mediating correlations. These contrasting signatures highlight a deeper asymmetry in the dynamical role of the mediator across the two regimes and may have implications for relativistic systems [18, 21].
Crucially, these results have significant implications for experimental tests of quantum gravity, such as BMV-type scenarios where two massive systems interact via a presumed quantum mediator. Our findings suggest that bipartite entanglement alone may offer only a partial diagnostic of the mediator’s quantum nature. Specifically, the absence of observed entanglement does not necessarily imply classical mediation. Instead, it may indicate that the mediator (analogous to oscillator B) is operating within a dynamically constrained subspace of its Hilbert space. This corresponds to the singular limit $\lambda_{B}\to\omega_{B}$ (HMR), wherein its ability to generate and transmit quantum correlations is effectively frozen, despite being fundamentally a quantum system. The observed vanishing entanglement in these specific parameter regions, even for non-zero quantum interactions, presents a critical caveat for entanglement-based protocols.
This singular limit, where the mediator’s kinetic energy is suppressed and its position becomes effectively static, may have a deeper physical interpretation within the context of quantum gravity. Given that gravity is inherently self-interacting and highly non-linear, such constrained dynamical behavior could be reminiscent of gravitational memory effects. In the context of BMV-type proposals, this suggests that even within a fully quantum framework of gravity, certain quasi-classical features (or strong self-interactions analogous to the self-squeezing term) could drive the mediating gravitational field into a subspace where it is dynamically inert and unable to coherently transmit quantum correlations. This implies that the quantum signature of gravity might be masked under specific, physically plausible conditions.
Therefore, for experimental tests aiming to definitively establish the quantum nature of gravity, it is paramount to account for such subtle dynamical constraints. Protocols must ensure that the hypothesized quantum mediator (e.g., gravitons or their effective field) operates in a regime where its quantum dynamical role is active and observable, rather than being effectively “frozen” into a quasi-classical configuration by strong self-interactions or specific environmental couplings. Our work provides a concrete, integrable model demonstrating how entanglement can vanish in a quantum system with non-zero interactions, serving as a critical benchmark for design and interpretation of future quantum gravity experiments.
@@ Appendix A Symplectic formalism
This section provides a brief overview of the covariance matrix and symplectic formalism, which is particularly well-suited for systems governed by Hamiltonian that are quadratic in the canonical quadrature operators ($\hat{q}$ and $\hat{p}$), or equivalently in the Bosonic creation and annihilation operators ($\hat{a}$ and $\hat{a}^{\dagger}$). Instead of working in the phase-space picture via wavefunctions or characteristic functions of type “$s$” (denoted $\chi_{s}$), a more elegant and computationally efficient approach involves the use of the covariance matrix formalism, especially for Gaussian states. For Gaussian states, knowledge of the first and second moments is sufficient to fully reconstruct the state, as all higher-order moments can be expressed in terms of them due to Wick’s theorem [50].
@@
of length “$2n$”, where “$n$” represents the number of oscillator modes in the system. The canonical commutation relations between the various mode operators can now be recast as
@@
In the traditional density matrix formalism the time evolution of the system is given by a unitary operator $\hat{U}(t)$ such that $\rho_{0}\rightarrow\rho(t)=\hat{U}(t)\rho_{0}\hat{U}^{{\dagger}}(t)$, which in the covariance matrix formalism corresponds to a symplectic transformation of the first and second order moments. The first order moments transform as $\langle\hat{\mathbb{X}}_{0}\rangle\rightarrow\langle\hat{\mathbb{X}}_{0}\rangle(t)=S_{H}(t)\langle\hat{\mathbb{X}}_{0}\rangle$ while the second order moments in the form of the covariance matrix transform as [51]:
@@
$A=\begin{pmatrix}(\bm{L}+\bm{O})+i(\bm{N}-\bm{M})&(\bm{L}-\bm{O})+i(\bm{N}+\bm{M})\\ (\bm{L}-\bm{O})-i(\bm{N}-\bm{M})&(\bm{L}+\bm{O})-i(\bm{N}-\bm{M})\end{pmatrix}$
Thus, any matrix in the complex basis would have the same structure as that of Eq (16). For the case where $U$ and $V$ are real matrices, $H$ becomes a real symmetric positive definite (the antisymmetric component of $H$ only contributes as a constant to the Hamiltonian and hence ignored  [52]) $2n\times 2n$ matrix. The symplectic time evolution matrix $S_{H}$ satisfies $S_{H}\Omega S_{H}^{{\dagger}}=S_{H}^{{\dagger}}\Omega S_{H}=\Omega$, and in general can be written as,
@@
Where the matrices $\bm{\alpha}$ and $\bm{\beta}$ are “$n\times n$”matrices called the Bogoliubov matrices. The symplectic condition $S_{0}\Omega S_{0}^{{\dagger}}=S_{0}^{{\dagger}}\Omega S_{0}=\Omega$ imposes further conditions on the Bogoliubov matrices called the Bogoliubov identities given as,
@@
Thus the entire system consists of solving the set of Eqs. (28) along with the constraint Eqs. (24), which gives the explicit expressions for $\bm{\alpha},\;\bm{\beta}$ and $\tilde{k}$. Once we obtain these quantities we can use Eqs. (27) to obtain $S_{H}(t)$ and then find $\langle\hat{\mathbb{X}_{0}}\rangle(t)$ and $\sigma(t)$. If at time “$t$”, the covariance matrix is given as ,
@@
The explicit form for $\bm{\mathcal{U}}(t)$ and $\bm{\mathcal{V}}(t)$ can be obtained using the explicit form of $S_{H}(t)$ from Eq (21), which gives us,

### v3

@@ IV Numerical Analysis and physical regimes
To move beyond the analytically tractable weak-coupling regime, we numerically simulate the system’s dynamics using the full Bogoliubov coefficients ($\bm{\alpha},\bm{\beta}$) (See Table I Appendix (B) for details). In what follows, we demonstrate our analysis on two distinct physical regimes starting from an initially squeezed state and provide physical arguments as to what is happening to the mediator in the two regimes
@@
    HMR (Heavy Mediator Regime): This regime corresponds to the limit $\lambda_{B}\to\omega_{B}$, with $1-\lambda_{A}/\omega_{A}$ remaining finite. In this limit, the mediator oscillator B behaves as an effectively heavy, quasi-classical field mode. To understand its physical implications, consider the single-oscillator Hamiltonian with self-squeezing (cf. Eq. (2a)). When expressed in the position–momentum basis, after transformation, the Hamiltonian simplifies to:
    $\hat{H}_{0B}={\hat{p}_{B}^{2}}/{(2\tilde{m}_{B}})+\tilde{k}_{B}\,\hat{x}_{B}^{2}/2,$ (9)
    Here, $\tilde{m}_{B}$ and $\tilde{k}_{B}$ are given by:
    $\tilde{m}_{B}=\frac{m_{B}}{1-\lambda_{B}/\omega_{B}},\tilde{\omega}_{B}^{2}=\omega_{B}^{2}-\lambda_{B}^{2},\tilde{k}_{B}=\tilde{m}_{B}\tilde{\omega}_{B}^{2},$ (10)
    where $m_{B}$ is the bare mass of the mediator. In the critical limiting case $\lambda_{B}\to\omega_{B}$, the effective mass $\tilde{m}_{B}$ diverges ($\tilde{m}_{B}\to\infty$) while the renormalized frequency vanishes ($\tilde{\omega}_{B}\to 0$). Crucially, their product $\tilde{k}_{B}=\tilde{m}_{B}\tilde{\omega}_{B}^{2}$ remains finite and equal to twice the bare spring constant $m_{B}\omega_{B}^{2}$. Physically, this causes the kinetic term to vanish in Hamiltonian (9), i.e.,
    $\hat{H}_{0B}\xrightarrow[\lambda_{B}\to\omega_{B}]{}\tilde{k}_{B}\,\hat{x}_{B}^{2}/2.$ (11)
    This limit corresponds to an extreme self-interaction that effectively suppresses the canonical momentum $\hat{p}_{B}$, making $\hat{x}_{B}$ a conserved quantity. Hence, the oscillator no longer evolves dynamically but acts as a static background field fixed by its initial position $\hat{x}_{B}(0)$. This behavior is analogous to a cosmological constant Peebles and Ratra [2003], where vacuum energy is given by the constant scalar field potential, here represented by a fixed energy density parameterized by $\hat{x}_{B}(0)^{2}$.
    Thus, in HMR, the mediator loses its ability to coherently store or transmit quantum information. Its Hilbert space remains intact, but its dynamical role is effectively frozen: the canonical momentum $\hat{p}_{B}$ vanishes from the Hamiltonian, and the position $\hat{x}_{B}$ becomes a conserved quantity. This reduction mirrors constrained dynamics in gauge theories and systems with Hamiltonian constraints, where certain degrees of freedom become dynamically redundant or decouple from the evolution Henneaux and Teitelboim [1992], Rovelli [2004].
@@
    LMR (Light Mediator Regime): This regime corresponds to $\lambda_{A}\to\omega_{A}$, while $1-\lambda_{B}/\omega_{B}$ remains finite. Here, mediator B is not significantly self-squeezed, remaining highly quantum and susceptible to fluctuations. Fig. (2) (bottom) contains the plot of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ as a function of $\omega_{B}$ and $g_{B}$ by fixing $\omega_{A}$, $\lambda_{A}$, $\lambda_{B}$, and $g_{A}$.
To visualize how entanglement between subsystems A and C depend on system parameters in both mediator regimes, we plot contour maps of the long-time averaged $\langle\mathcal{E}_{\mathcal{N}}\rangle$. For the HMR [Fig. 2 (top)], we take $(\omega_{B},\lambda_{B},\lambda_{A},g_{B})=(5.0,4.9,10^{-4},1.0)$ with initial state $r=(0.1,1.0,0.1)$ and $\theta=(-\pi/3,\pi/4,\pi/6)$, and plot $\langle\mathcal{E}_{\mathcal{N}}\rangle$ as a function of $(\omega_{A},g_{A})$. For the LMR [Fig. 2 (bottom)], we set $(\omega_{A},\lambda_{A},\lambda_{B},g_{B})=(5.0,4.999,\omega_{B}/1.5,1.4)$ with $r=(0.5,2.0,0.5)$ and $\theta=(0,0,0)$ and plot $\langle\mathcal{E}_{\mathcal{N}}\rangle$ as a function of $(\omega_{B},g_{A})$. From numerical simulations, we also confirm that similar behavior is obtained even for the case where the system starts out in a vacuum state initially. Further $\langle{\cal D}\rangle$ shows similar behaviour as that of entanglement (see Figs. 4 and 5, Appendix (E)) . The contour plots of $\langle\mathcal{E}_{\mathcal{N}}\rangle$ in both the LMR and HMR regimes indicate that entanglement between the two subsystems can be mediated via oscillator $B$, though through qualitatively distinct mechanisms. To further differentiate these regimes and characterize the mediator’s behavior, we compute the time-averaged dynamical fidelity susceptibility $\langle\chi_{F}\rangle$ Chandran and Shankaranarayanan [2024]. In order to capture the mediator’s effect on the two subsystems, this quantity is evaluated for the reduced state of AACC, described by the covariance matrix $\sigma_{AC}(t)$. The dynamical fidelity $\mathcal{F}(t)$ for mixed states is given by
$\mathcal{F}_{\text{mixed}}=\mathrm{Tr}\!\left[\sqrt{\rho^{1/2}(t)\,\rho(t+\delta t)\,\rho^{1/2}(t)}\right].$ (12)
Expanding $\mathcal{F}(t)$ for small $\delta t$ yields
$\!\!\!\!\mathcal{F}(t,\delta t)\sim 1-\chi_{F}(t)\,(\delta t)^{2},$~ $~\chi_{F}(t)\equiv-\tfrac{1}{2}\,\left.\frac{d^{2}\mathcal{F}}{d(\delta t)^{2}}\right|_{\delta t=0}$ (13)
where $\chi_{F}(t)$ is the dynamical fidelity susceptibility (DFS). To compute $\chi_{F}(t_{0})$, we take $\sigma_{AC}(t_{0})$ as the reference state and evaluate its overlap with $\sigma_{AC}(t_{0}\pm\delta t)$ via the Gaussian fidelity expression provided in Appendix (C), from which $\chi_{F}$ is extracted. We then evaluate the time-averaged $\langle\chi_{F}\rangle$ as a function of the mediator frequency $\omega_{B}$, as shown in Fig. (3). Our results reveal distinct variations in $\langle\chi_{F}\rangle$ depending on the system’s proximity to the HMR. For parameter values ($\lambda_{B}$) away from the HMR, the susceptibility exhibits a characteristic inflection point in the vicinity of $\omega_{B}\approx\omega_{A}$ — a feature absent in the near-HMR and exact HMR regimes. In this regime, $\langle\chi_{F}\rangle$ shows heightened sensitivity only within a narrow region around $\omega_{B}\approx\omega_{A}$, saturating to different baseline values elsewhere. By contrast, in the proximity of the HMR limit ($\lambda_{B}\approx\omega_{A}$), the response of $\langle\chi_{F}\rangle$ sharpens into a prominent, resonance-like peak. Beyond this peak, the susceptibility rapidly falls off and saturates to the flat HMR value corresponding to the critical point $\lambda_{B}=\omega_{B}$. The nature of this sensitivity fundamentally differs across regimes: near the HMR, $\langle\chi_{F}\rangle$ is highly sensitive to variations in $\omega_{B}$ away from $\omega_{A}$, reflecting a broad change in the mediator’s dynamics. Conversely, far from the HMR, the susceptibility is enhanced only in the immediate vicinity of $\omega_{B}\approx\omega_{A}$, indicating a localized sensitivity. These distinct trends powerfully demonstrate that $\langle\chi_{F}\rangle$ serves as a robust diagnostic of the mediator’s operational regime. Plots for the case $\lambda_{A}=0.01\,\omega_{A}$ are provided in Appendix (G).
In our model, the mediator’s intrinsic character is determined by the parameters $\omega_{B}$ and $\lambda_{B}$. Varying $\omega_{B}$ effectively simulates different mediator regimes. Within BMV-type experimental scenarios, where the mediator is gravitational, tuning effective parameters analogous to $\omega_{B}$ could correspond to probing different gravitational regimes, such as variations in the local spacetime curvature $\mathcal{R}$. The quantity $\sqrt{\mathcal{R}}$ then sets a natural frequency scale associated with the hypothesized gravitational mediator. In such a setting, system parameters such as $\omega_{A}$ can be chosen to be comparable to $\sqrt{\mathcal{R}}$, and varied around this scale to explore the behavior of the time-averaged dynamical fidelity susceptibility. Measuring and comparing $\langle\chi_{F}\rangle$ across such regimes could provide operational insight into whether the mediator exhibits fundamentally quantum or effectively classical behavior—thereby offering a novel experimental pathway to characterize gravitationally-mediated interactions.
Figure 3: Normalized time-averaged $\langle\chi_{F}\rangle$ vs $\omega_{B}$ for two different initial states (vacuum: left, squeezed: right) with $r_{1}=r_{3}=0.5$, $r_{2}=0.2$, $\theta_{1}=\theta_{2}=\theta_{3}=0$. Fixed parameters: $\omega_{A}=5$, $\lambda_{A}=0.5\,\omega_{A}$, $g_{A}=g_{B}=0.6$.
@@ V Results and Implications
Our numerical analysis, presented in Fig. (2), reveals that in both the HMR and the LMR, the mediator can facilitate the establishment of quantum correlations between A and C, but with different structural and spectral features in the two regimes. The LMR consistently displays a smoother and more continuous entanglement landscape, with nonzero quantum correlations broadly distributed across wide parameter ranges [Fig. 2 (bottom), Fig. 4(c) in Appendix (E)]. This indicates a robust ability of the dynamically active, “light” mediator $B$ to facilitate entanglement between $A$ and $C$. When the mediator can fluctuate significantly, it readily mediates quantum correlations. While at isolated instants the bipartite entanglement between $A$ and $C$ may transiently vanish due to the global state resembling a continuous variable-GHZ state, our analysis captures the time-averaged behavior over dynamical cycles, which always yields nonzero correlations in this regime.
In contrast, in the HMR, bipartite entanglement emerges only within a sharply localized patches of the $(\omega_{A},g_{A})$ parameter space [Fig. 2 (top), Fig. 4(a) in Appendix (E)], while outside this region the time-averaged logarithmic negativity is strongly suppressed or vanishes in the exact limit ($\lambda_{B}=\omega_{B}$)(See Appendix (G) for details). This arises from fine-tuned parameter choices that bring the system close to resonance, enabling transient enhancement even at finite coupling strengths $g_{A}$ and $g_{B}$. By contrast, in the LMR such suppression only occurs for very small coupling strengths.
This behaviour is reminiscent of quantum many-body scars, in which non-generic eigenstates maintain unexpectedly low entanglement and resist thermalization Moudgalya et al. [2022], Deutsch [2018]. Here, we observe an analogous form of scars, not in a many-body Hilbert space but within the space of Gaussian states of this continuous-variable system. Further, the suppression of entanglement observed in the HMR may superficially resemble entanglement sudden death, but in our integrable Gaussian setting complete disappearance of correlations occurs only in the strict zero-mode limit of the mediator (for details, see Appendix (F)).
This asymmetry arises from the eigenmode structure: in HMR ($\lambda_{B}\to\omega_{B}$), the mediator’s near degenerate eigenmodes ($k_{2}\to 0$, $k_{1}\to k_{3}$) induce entanglement suppression, demonstrating that quantum systems can appear classical under dynamical constraints, while in LMR ($\lambda_{A}\to\omega_{A}$), the two eigenmodes vanish simultaneously ($k_{1},k_{2}\to 0$), with the third remaining finite. This suggests persistent entanglement arises from the mediator’s active fluctuations confirming its quantum role. These contrasting signatures highlight an inherent asymmetry in the mediator’s dynamical role across the two regimes and may have implications for relativistic systems Fragkos et al. [2022], Martín-Martínez and Perche [2023].
These results carry significant implications for experimental tests of quantum gravity, such as BMV-type scenarios where two massive systems interact via a presumed quantum mediator. Such scenarios are being explored with optomechanical oscillators and trapped ions Teufel et al. [2011], Chan et al. [2011], Abbott et al. [2009], Krisnanda et al. [2020]. Our findings demonstrate that a mediator treated as quantum can establish entanglement while existing in two physically distinct regimes: in the LMR it actively establishes entanglement, while in the HMR it becomes dynamically constrained and correlations vanish at ($\lambda_{B}=\omega_{B}$). Thus, Entanglement-based witnesses alone may therefore be insufficient, as they cannot distinguish these regimes by themselves. Thus, absence of entanglement in BMV-type tests may indicate HMR-like freezing not classicality.
When extended to gravity, our results show that a quantum mediator can exhibit signatures that appear classical, namely suppressed entanglement and frozen fluctuations. The dynamical fidelity susceptibility ($\langle\chi_{F}\rangle$) distinguishes these regimes, revealing when the mediator is dynamically constrained yet still quantum. Such constrained behavior may carry deeper meaning in quantum gravity, where self-interactions can localize information in the field, reminiscent of gravitational memory effects. Consequently, expected quantum signatures may be masked under realistic conditions, underscoring the need for probes beyond entanglement.
For BMV-type experiments, this means the mediator must operate in a regime where its quantum dynamics remain active—a nontrivial requirement given our limited knowledge of its internal structure. We show that tracking ($\langle\chi_{F}\rangle$) against system parameters, such as oscillator frequencies or effective mass, provides a direct handle on the predicted HMR–LMR transition. Thus, the BMV paradigm must be broadened: entanglement generation is sufficient but not necessary to establish a quantum mediator. Revised criteria should combine entanglement with complementary probes such as mediator tomography Vidal et al. [2025] and fidelity susceptibility.
@@ Appendix A Symplectic formalism
This section provides a brief overview of the covariance matrix and symplectic formalism, which is particularly well-suited for systems governed by Hamiltonian that are quadratic in the canonical quadrature operators ($\hat{q}$ and $\hat{p}$), or equivalently in the Bosonic creation and annihilation operators ($\hat{a}$ and $\hat{a}^{\dagger}$). Instead of working in the phase-space picture via wavefunctions or characteristic functions of type “ss” (denoted $\chi_{s}$), a more elegant and computationally efficient approach involves the use of the covariance matrix formalism, especially for Gaussian states. For Gaussian states, knowledge of the first and second moments is sufficient to fully reconstruct the state, as all higher-order moments can be expressed in terms of them due to Wick’s theorem Weedbrook et al. [2012].
@@
of length “2n2n”, where “nn” represents the number of oscillator modes in the system. The canonical commutation relations between the various mode operators can now be recast as
@@
In the traditional density matrix formalism the time evolution of the system is given by a unitary operator $\hat{U}(t)$ such that $\rho_{0}\rightarrow\rho(t)=\hat{U}(t)\rho_{0}\hat{U}^{{\dagger}}(t)$, which in the covariance matrix formalism corresponds to a symplectic transformation of the first and second order moments. The first order moments transform as $\langle\hat{\mathbb{X}}_{0}\rangle\rightarrow\langle\hat{\mathbb{X}}_{0}\rangle(t)=S_{H}(t)\langle\hat{\mathbb{X}}_{0}\rangle$ while the second order moments in the form of the covariance matrix transform as Simon et al. [1988]:
@@
$A=\begin{pmatrix}(\bm{L}+\bm{O})+i(\bm{N}-\bm{M})&(\bm{L}-\bm{O})+i(\bm{N}+\bm{M})\\ (\bm{L}-\bm{O})-i(\bm{N}-\bm{M})&(\bm{L}+\bm{O})-i(\bm{N}-\bm{M})\end{pmatrix}$ (22)
Thus, any matrix in the complex basis would have the same structure as that of Eq. (17). For the case where $U$ and $V$ are real matrices, $H$ becomes a real symmetric positive definite (the antisymmetric component of $H$ only contributes as a constant to the Hamiltonian and hence ignored  Serafini [2017]) $2n\times 2n$ matrix. The symplectic time evolution matrix $S_{H}$ satisfies $S_{H}\Omega S_{H}^{{\dagger}}=S_{H}^{{\dagger}}\Omega S_{H}=\Omega$, and in general can be written as,
@@
Where the matrices $\bm{\alpha}$ and $\bm{\beta}$ are $n\times n$”matrices called the Bogoliubov matrices. The symplectic condition $S_{0}\Omega S_{0}^{{\dagger}}=S_{0}^{{\dagger}}\Omega S_{0}=\Omega$ imposes further conditions on the Bogoliubov matrices called the Bogoliubov identities given as,
@@
Thus the entire system consists of solving the set of Eqs. (30) along with the constraint Eqs. (26), which gives the explicit expressions for $\bm{\alpha},\;\bm{\beta}$ and $\tilde{k}$. Once we obtain these quantities we can use Eqs. (29) to obtain $S_{H}(t)$ and then find $\langle\hat{\mathbb{X}_{0}}\rangle(t)$ and $\sigma(t)$. If at time “tt”, the covariance matrix is given as ,
@@
The explicit form for $\bm{\mathcal{U}}(t)$ and $\bm{\mathcal{V}}(t)$ can be obtained using the explicit form of $S_{H}(t)$ from Eq. (23), which gives us,